# Emergence of Goal-Directed Behaviors via Active Inference with Self-Prior

### v1

@@ 4.2 Continuous Simulation
We conducted the continuous environment experiments on a machine running Ubuntu 22.04.5 LTS 64-bit (Linux 5.15.0-1066) with an Intel Xeon E5-2698 v4 CPU and an NVIDIA Tesla V100-SXM2 GPU, using Python 3.11.10 and PyTorch 2.5.1. For our normalizing-flows-based self-prior, we employed the “zuko” library (version 1.3.0), and we based much of our overall algorithmic implementation on Dreamer (Hafner et al.,, 2020) and Contrastive Active Inference (Mazzaglia et al.,, 2021).
To train and evaluate our deep model, we stored episodes in a replay buffer and randomly sampled them during model training. Each episode was collected either by using random actions or by following the agent’s policy from active inference, chosen with equal probability (50%). Likewise, for half of the episodes, a sticker was placed at a random location on the arm; for the other half, no sticker was used. Each episode consisted of $L=1000$ time steps.
After collecting an initial batch of 10 episodes, we trained the model for 100 epochs at the end of each episode. In each epoch, we sampled $B=50$ trajectories of length $L=50$ from the replay buffer, and generated imagined rollouts with a planning horizon $H=15$ for policy learning. For variational free energy minimization, trajectories were sampled from all available data—regardless of the presence of a sticker or whether the behavior was random or policy-driven. This was intended to ensure the acquisition of an accurate world model, similar to the assumption in the discrete environment that $\mathbf{A}$ and $\mathbf{B}$ were already known. Likewise, all data were used for expected free energy minimization, so that the agent could learn which actions reduce free energy effectively across various conditions.
@@
Figure 8: Agent behavior when a sticker is placed in the continuous environment. (a) When a sticker (blue circle) is attached, the hand (red circle) moves toward it, illustrating goal-directed reaching behavior. The figure visualizes the agent’s tactile matrix, where grayscale tactile data is overlaid with colored markers for the hand and sticker for clarity. (b) Reaching for the sticker reduces expected free energy, and removing the sticker leads to its minimization. Shaded areas represent the standard deviation across 64 experiments using 8 model training seeds tested on 8 environment seeds. The figure uses seed 14, selected for clearly demonstrating the reaching behavior.
The agent displayed goal-directed reaching behavior whenever a sticker was placed on the left arm (Figure 8(a)). This occurred because touching the sticker minimized free energy by aligning with the self-prior (Figure 8(b)). Moreover, after successfully reaching the sticker, the agent persisted in touching it until the sticker was removed, indicating a consistent drive to resolve mismatch rather than merely losing interest once contact was made.
Interestingly, the agent exhibited reaching behavior toward the sticker even when the right hand was not initially touching the arm. If the self-prior had been learned based solely on tactile information, (i) the case where the left arm is touched by the right hand and (ii) the case where a sticker is attached might not be distinguishable—since both yield a single tactile point on the arm. In that case, no reaching behavior toward the sticker would emerge. However, the agent did reach toward the sticker, suggesting that the self-prior was learned by integrating both tactile and proprioceptive inputs. In this sense, our system, which utilizes a self-prior learned from multimodal sensory inputs to generate reaching behavior, resembles the concept of a body schema mentioned in (Gallagher,, 1986; Hoffmann,, 2021), which supports action planning and control.
Additionally, even if extra touches arose mid-trajectory—e.g., from the right hand itself—thus momentarily increasing mismatch relative to a single-point self-prior, the agent accepted this short-term deviation to achieve the ultimate goal of minimizing overall free energy. This robust, noise-tolerant sequence of goal-directed action demonstrates that our agent maintained a coherent intention throughout its interactions.
@@ 5 Discussion
We proposed the self-prior as an intrinsic preference that is autonomously formed from an agent’s sensory experiences. We then applied it to a simulated infant agent within the active inference framework and confirmed the emergence of goal-directed behavior—specifically, reaching for a sticker. In this section, to clarify the contributions of our work to developmental science, we first provide a detailed discussion of how the self-prior can be classified within existing computational models of intrinsic motivation. We then propose the role of the self-prior as a potential origin of early intentional behavior. After that, we explain how our approach differs from previous studies that attempted to generate spontaneous behaviors, and finally, we discuss the limitations of this work.
5.1 Internal Homeostatic Intrinsic Motivation
The self-prior represents an intrinsic motivation term that has been separated from extrinsic motivation, which may seem counterintuitive. Furthermore, the idea that intrinsic motivation can emerge from sensory data originating from the external world might appear paradoxical. We argue that this confusion stems from the inconsistent use of the terms “external” and “internal,” as also noted by (Oudeyer & Kaplan,, 2009). In this paper, we adopt the classification they recommend, distinguishing between:
  • External vs. internal, referring to whether the reward computation occurs in the environment or within the agent.
  • Extrinsic vs. intrinsic, referring to whether the criterion for reward is determined externally or internally.
According to these definitions, any external reward—such as a game score in machine learning tasks—qualifies as extrinsic, because it is provided from outside the agent, along with its preset threshold. Meanwhile, an internal drive can be either extrinsic or intrinsic. An internal drive is considered intrinsic if the agent calculates its own reward threshold based on its internal model—such as measuring “new information gain” according to a learned world model. By contrast, an internal yet extrinsic motivation, which can be confusing at first, is one in which the agent checks its own state and directly computes a reward based on a criterion that is externally fixed, for example “maintaining a battery level of 50%.”
Thus, the decisive factor for distinguishing extrinsic versus intrinsic motivation is not whether the information originates in the environment or within the agent, but rather who sets the threshold for reward computation. In this sense, the expected free energy remains an internal framework, since the agent itself computes both extrinsic and intrinsic components. Moreover, the self-prior approach introduced here qualifies as internal intrinsic motivation because the agent’s internal model (the self-prior) directly calculates and updates the criterion for behavior.
(Oudeyer & Kaplan,, 2009) also introduce another classification axis—homeostatic versus heterostatic—and note that homeostatic intrinsic motivation can indeed exist. In other words, if the agent’s drive is directed toward acquiring new information, it constitutes a heterostatic form of intrinsic motivation; however, maintaining a familiar state can also be considered an intrinsic motivation (homeostatic), because the agent internally defines the “familiarity” setpoint.
The expected free energy framework typically interprets the preferred prior as covering observations tied to survival requirements (e.g., blood sugar or battery level). Because such setpoints must not be altered arbitrarily, they are often considered extrinsic and not explicitly distinguished from their homeostatic nature. However, if the agent can freely specify its own setpoints, we can classify those setpoints as homeostatic intrinsic motivations.
Accordingly, we regard the motivation induced by the self-prior as an internal homeostatic intrinsic motivation—it aims to “maintain or restore a reference pattern” learned from the agent’s own sensory experiences (Table 2). This stands in contrast to the more traditional, heterostatic intrinsic motivations that seek new information for learning (e.g., reducing uncertainty about the hidden state). Instead, the self-prior directly triggers action to resolve any mismatch from the already acquired pattern of one’s own body. In (Oudeyer & Kaplan,, 2009), this is akin to distributional familiarity motivation (DFM), which drives the agent to return to sensory configurations it has repeatedly encountered.
Extrinsic Intrinsic
Homeostatic Heterostatic
External Score in video game (N/A) (N/A)
Internal Gap of sensory data from fixed value Gap of sensory data from self-prior Information gain by sensory data
@@ 5.2 Origin of Early Intentional Behaviors
Infants initially depend heavily on reflex mechanisms for action. Over time, however, they gradually develop what (Mele & Moser,, 1994) describes as “intentional behavior,” comprising (i) a motivational system that judges how to handle incoming stimuli and (ii) the capability to execute genuinely goal-directed actions. According to (Zaadnoordijk & Bayne,, 2020), stimulus-driven intentions arise first, and then, with accumulating experience, more endogenous intentions emerge. For instance, newborns rely on reflexes such as rooting to acquire feeding experience; later, merely seeing a bottle prompts them to reach out and drink (stimulus-driven intention). As they become aware of internal states like hunger, they begin actively searching for the bottle themselves (endogenous intention).
@@
Over time, to recognize themselves as independent entities, infants must undergo abundant experiences featuring multimodal redundancy and temporal–spatial contingencies (Rochat,, 1998). Sequentially developing intentional actions may spontaneously provide opportunities for such experiences. (Thelen et al.,, 1993) similarly noted that infants initially lack both the ability and the intention to perform reaching, and that the formation of stable intentions, coupled with active exploration driven by those intentions, facilitates the learning of reaching. However, prior studies have not examined in detail the neural architecture that underlies both stimulus-driven intentions and intentional actions.
@@
Meanwhile, neurophysiological evidence suggests that the anterior cingulate cortex primarily deals with reward-related errors in individual decisions, whereas the posterior cingulate cortex processes discrepancy signals related to autobiographical information—such as one’s own body or past experiences—and triggers new behaviors (Pearson et al.,, 2011; Brewer et al.,, 2013). Our proposed self-prior aligns with this notion of “self-relevant predictions,” indicating that when new input fails to satisfy these predictions, the system initiates action. In other words, even without any externally designed goal, an intrinsic drive to preserve a self-prior learned from past experiences can manifest behaviors akin to an infant persistently touching a newly discovered sticker on their arm. This mirrors Piaget’s concept of primary circular reactions, where repeated attempts can refine intentions further, offering a theoretical foundation for computational studies on the origins of infant intentional behavior.
@@
In our experiments, we fixed the self-prior’s learning rate. However, an excessively high rate could lead to overreactions (compulsive-like responses) to minor changes, whereas an overly low rate might cause indifference (apathy) toward new stimuli. Future work could investigate how more flexible learning rates—modeled after biological variability—shape different behavioral styles via self-relevant prediction-error processing.
5.3 Computation models for Intrinsic Motivation
Approaches that aim to computationally address intrinsic motivation have primarily been explored in reinforcement learning, where designing high-quality reward functions remains a significant challenge. As a result, intrinsic motivation has been proposed as a means for agents to more efficiently explore the environment and better infer reward functions. However, a fundamental question persists: how can an agent autonomously form motivation and generate behavior in environments where no goals are given, and what kind of value guides the agent’s actions under such conditions (Juechems & Summerfield,, 2019)?
An alternative approach is homeostatic reinforcement learning, in which the agent can generate a range of behaviors simply by maintaining certain sensory channels to specified thresholds, without environmental goal (Keramati & Gutkin,, 2011). Recent studies have extended this concept to complex settings using deep neural networks, examining how internal physiological states drive behavior (Yoshida et al.,, 2024). Under active inference, a similar mechanism emerges: by minimizing free energy with respect to an internal setpoint (expressed as a preferred prior), an agent can generate a single goal-directed act like reaching for a target (Oliver et al.,, 2022) or orchestrate repeated behaviors (Matsumoto et al.,, 2022).
However, although such methods may appear to operate without external rewards, most rely on the experimenter fixing the reference values (e.g., setpoints, preferred priors) in advance. Thus, while the agent’s reward calculation is internal, the ultimate source of its setpoint is extrinsic. As a result, these systems do not fully resolve the fundamental question of “Where does value come from, and how can the agent establish it autonomously?”
To allow the agent to establish its own value function independently, neither external rewards nor internal setpoints can be predetermined. But without these metrics, it is difficult to track progress or judge the utility of behaviors, which is why many studies on intrinsic motivation still use it primarily to facilitate exploration for extrinsic tasks (Aubret et al.,, 2023)—with limited focus on how the agent might self-generate and maintain goals.
Some works do investigate the behaviors arising purely from intrinsic motivation and qualitatively evaluate them in scenarios with no externally defined reward (Eysenbach et al.,, 2018; Pathak et al.,, 2017). While these highlight the agent’s acquisition of diverse skills or behaviors, they typically involve a heterostatic approach—continuously seeking novel states. Consequently, once uncertainty is sufficiently reduced, further exploration tends to diminish, and fewer new behaviors emerge.
Active inference offers a unified theoretical framework that can integrate both extrinsic and intrinsic motivations under the single quantitative measure of free energy (Parr et al.,, 2022). Furthermore, (Biehl et al.,, 2018) showed that many existing methods for intrinsic motivation can be cast within active inference as processes for improving the posterior—a perspective reminiscent of (Schmidhuber,, 2010), who defined intrinsic motivation in terms of seeking “better world models.” However, such posterior-improving motives are likewise heterostatic: they promote exploration until uncertainty is resolved, after which they elicit fewer notable actions.
Recently, studies have begun to explore homeostatic forms of intrinsic motivation, in which agents autonomously define their own setpoints but aim to preserve already learned states or past experiences, rather than continuously seeking novelty. For example, (Marcel et al.,, 2022) proposed a model in which a latent representation of self-touch is first learned and then used as a setpoint to guide the current observation toward the closest self-touch state, thereby inducing reaching behavior. Similarly, (Takemura et al.,, 2018; Kim et al.,, 2023) treated one-step forward and inverse models of a robot’s end-effector—learned from motor babbling—as behavioral setpoints, and demonstrated that such models could generate reaching toward distant targets without explicit rewards. Closer to our approach, (Sajid et al.,, 2021) introduced the concept of "learned preferences" within the free energy framework and semantically analyzed how a drive to reduce mismatches between observation and preference can lead to new behaviors. However, these approaches do not sufficiently discuss how such self-acquired setpoints could be integrated with conventional extrinsic or information-gain-based intrinsic motivations.
In this paper, we address these questions within the active inference framework by proposing a method whereby the agent shapes its own setpoint based on past experience, then spontaneously generates goal-directed behavior whenever incoming sensations deviate from this learned internal model. Our approach answers the question “Where does value come from?” by suggesting that “part of value may be shaped by the agent’s accumulated experience,” demonstrating that (i) the agent maintains self-defined setpoints instead of adhering to experimenter-imposed ones; (ii) it still utilizes the active inference machinery for long-horizon planning to minimize expected free energy; and (iii) it can be theoretically integrated with existing extrinsic or heterostatic intrinsic motivations. The result is a comprehensive framework wherein the agent’s value system is shaped and maintained internally, yet remains fully compatible with established motivational processes.
@@ 5.4 Limitations and Future Directions
Although our study provides a proof of concept, it is subject to several limitations. First, we tested our approach in a discrete environment and a low-degree-of-freedom (3-DoF) continuous environment. Real infants or robots, however, exhibit far more complex and varied sensorimotor interactions, underscoring the need to extend our framework to higher-dimensional state and action spaces. In particular, future work could draw on simulations of early human sensorimotor experience (including fetal stages) (Kim et al.,, 2022) to explore how rich, in-utero sensorimotor inputs shape self-priors and goal-directed actions, clarifying the types of actions that might emerge under embodied developmental constraints (Kuniyoshi,, 2019).
Additionally, our experiments focused on scenarios wherein the sticker is introduced abruptly. This choice helped isolate the effects of familiarity-based intrinsic motivation by excluding a fixed extrinsic motivation (such as externally imposed reward criteria). In reality, infants and robots must often balance multiple drives (e.g., energy homeostasis and novelty-driven information gain). Designing experiments that systematically assess how different motivations complement or conflict—and how these are integrated into coherent behavior—remains an important avenue for future research.
@@
Our study demonstrated that, by separating an intrinsic homeostatic term based on a learnable self-prior from the extrinsic term based on a fixed preferred prior—commonly used in prior work with expected free energy—an agent can generate not only actions aimed at achieving predefined goals but also spontaneously generate goal-directed behaviors. However, a current limitation is that while the use of the self-prior is integrated into policy generation via expected free energy, the learning of the self-prior is performed separately from the learning of variational free energy.
@@
One potential clue to resolving this separation lies in a hierarchical architecture. Specifically, a higher-level hidden state that reflects long-term experiences may be gradually learned, and the likelihood distribution inferred from it could serve as the self-prior for the lower level. This differs from conventional hierarchical inference, where the higher level simply learns priors and determines actions on a slower timescale to influence the lower level (e.g., $P(s_{t}|s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$). Instead, it continuously receives information from the lower level to learn a prior, serving only to form a self-prior over a long duration (e.g., $P(s)=\text{Cat}(\mathbf{B})$).
@@
In the discrete environment, we saw that placing a sticker and leaving it in place gradually altered the self-prior and, consequently, the agent’s behavior. By contrast, in the continuous environment, we fixed the probability of encountering a sticker at 5%, effectively decoupling self-prior learning from the evolving incidence of sticker observations. This occurred because the self-prior model learned from a replay buffer that stored past data without considering time in sequence, making it difficult to investigate incremental changes in the self-prior over time. To model how a stimulus-driven intention might become consolidated into a more endogenous form of intention in real infants or robots, we may need hierarchical state-space models capable of long-range temporal dependencies (Friston et al.,, 2017) or transformer-based architectures (Vaswani et al.,, 2017) that can learn world models from extended temporal sequences (Chen et al.,, 2022).

### v2

@@ 4.2 Simulation in Continuous Environment
We conducted the continuous environment experiments on a machine running Ubuntu 22.04.5 LTS 64-bit (Linux 5.15.0-1066) with an Intel Xeon E5-2698 v4 CPU and an NVIDIA Tesla V100-SXM2 GPU, using Python 3.11.10 and PyTorch 2.5.1. For our normalizing flows based self-prior, we employed the “zuko” library (version 1.3.0), and we based much of our overall algorithmic implementation on Dreamer (hafner_dreamer_2020) and Contrastive Active Inference (mazzaglia_contrastive_2021).
To train and evaluate our deep model, we stored episodes in a replay buffer and randomly sampled them during model training. Each episode was collected either by using random actions or by following the agent’s policy from active inference, chosen with equal probability (50%). Likewise, for half of the episodes, a sticker was placed at a random location on the arm; for the other half, no sticker was used.
After collecting an initial batch of 100 episodes, we trained the model for 100 epochs at the end of each episode. In each epoch, we sampled $B=50$ trajectories of length $L=50$ from the replay buffer, and generated imagined rollouts with a planning horizon $H=15$ for policy learning. For variational free energy minimization, trajectories were sampled from all available data, regardless of the presence of a sticker or whether the behavior was random or policy-driven. This was intended to ensure the acquisition of an accurate world model, similar to the assumption in the discrete environment that $\mathbf{A}$ and $\mathbf{B}$ were already known. Likewise, all data were used for expected free energy minimization, so that the agent could learn which actions reduce free energy effectively across various conditions.
@@
After collecting and training for 2,000 episodes, we attached a sticker to the agent’s left arm. As a result, the agent exhibited goal-directed behavior of reaching toward the sticker (Figure 8(a)). This occurred because touching the sticker aligned with the self-prior, minimizing free energy (Figure 8(b)). Moreover, after reaching the sticker, the agent continued touching it and eventually removed it, indicating that it had a persistent motivation to reduce the mismatch rather than simply losing interest after contact.
@@
Figure 8: Agent behavior when a sticker is placed in the continuous environment. (a) When a sticker (blue circle) is attached, the hand (red circle) moves toward it, illustrating goal-directed reaching behavior. The figure visualizes the agent’s tactile matrix, where grayscale tactile data is overlaid with colored markers for the hand and sticker for clarity. (b) Reaching for the sticker reduces expected free energy, and removing the sticker leads to its minimization. Shaded areas represent the standard deviation across 64 experiments using 8 model training seeds (0˜7) tested on 8 environment seeds (0˜7). The figure uses seed 4, selected for clearly demonstrating the reaching behavior.
Interestingly, the agent exhibited reaching behavior toward the sticker even when the right hand was not initially touching the arm. If the self-prior had been learned based solely on tactile information, (i) the case where the left arm is touched by the right hand and (ii) the case where a sticker is attached might not be distinguishable, since both yield a single tactile point on the arm. In that case, no reaching behavior toward the sticker would emerge. However, the agent did reach toward the sticker, suggesting that the self-prior was learned by integrating both tactile and proprioceptive inputs. In this sense, our system, which utilizes a self-prior learned from multimodal sensory inputs to generate reaching behavior, resembles the concept of a body schema mentioned in (gallagher_body_1986; hoffmann_body_2021), which supports action planning and control.
Additionally, even if extra touches arose mid-trajectory (e.g., from the right hand itself) thus momentarily increasing mismatch relative to a single-point self-prior, the agent accepted this short-term deviation to achieve the ultimate goal of minimizing overall free energy. This robust, noise-tolerant sequence of goal-directed action demonstrates that our agent maintained a coherent intention throughout its interactions.
@@ 5 Discussion
We proposed the self-prior as an intrinsic preference that is autonomously formed from an agent’s sensory experiences. We then applied it to a simulated infant agent within the active inference framework and confirmed the emergence of goal-directed behavior, specifically reaching for a sticker. In this section, to clarify the contributions of our work to developmental science, we explain how our approach differs from previous studies that attempted to generate spontaneous behaviors, propose the role of the self-prior as a potential origin of early intentional behavior, and finally discuss the limitations of this work and future research directions.
5.1 Computation Models of Intrinsic Motivation
Approaches that aim to computationally address intrinsic motivation have primarily been explored in reinforcement learning, where designing high-quality reward functions remains a significant challenge. As a result, intrinsic motivation has been proposed as a means for agents to more efficiently explore the environment and better infer reward functions. However, a fundamental question persists: how can an agent autonomously form motivation and generate behavior in environments where no goals are given, and what kind of value guides the agent’s actions under such conditions (juechems_where_2019)?
Extrinsic Homeostatic Reinforcement Learning
An alternative approach is homeostatic reinforcement learning, in which the agent can generate a range of behaviors simply by maintaining certain sensory channels to specified thresholds, without environmental goal (keramati_reinforcement_2011). Recent studies have extended this concept to complex settings using deep neural networks, examining how internal physiological states drive behavior (yoshida_emergence_2024). Under active inference, a similar mechanism emerges: by minimizing free energy with respect to an internal setpoint (expressed as a preferred prior), an agent can generate a single goal-directed act like reaching for a target (oliver_empirical_2022) or orchestrate repeated behaviors (matsumoto_goal-directed_2022).
However, although these approaches appear intrinsic because behaviors occur without external rewards, most have a limitation in that the experimenter fixes the reference values (e.g., setpoints or preferred priors) in advance. That is, even though the agent calculates rewards internally, the reference itself is externally provided, so it ultimately has an extrinsic nature. Therefore, these approaches do not fully resolve the fundamental question of "Where does value come from, and can the agent establish it by itself?"
Reward-auxiliary Intrinsic Motivation
To allow the agent to establish its own value function independently, neither external rewards nor internal setpoints can be predetermined. But without these metrics, it is difficult to track learning progress or judge the utility of behaviors, which is why many studies on intrinsic motivation still tend to analyze it primarily as a means to facilitate exploration for extrinsic tasks (aubret_information-theoretic_2023). That is, the process by which agents generate and maintain goals themselves has been relatively less illuminated.
Heterostatic Intrinsic Motivation
Some works do investigate the behaviors arising purely from intrinsic motivation and qualitatively evaluate them in scenarios with no externally defined reward (eysenbach_diayn_2018; pathak_icm_2017). While these highlight the agent’s acquisition of diverse skills or behaviors, they typically involve a heterostatic approach, continuously seeking novel states. Consequently, once uncertainty is sufficiently reduced, further exploration tends to diminish, and fewer new behaviors emerge.
Active inference offers a unified theoretical framework that can integrate both extrinsic and intrinsic motivations under the single quantitative measure of free energy (parr_active_2022). Furthermore, (biehl_expanding_2018) showed that many existing methods for intrinsic motivation can be cast within active inference as processes for improving the posterior, a perspective reminiscent of (schmidhuber_formal_2010), who defined intrinsic motivation in terms of seeking “better world models.” However, such posterior-improving motives are likewise heterostatic: they promote exploration until uncertainty is resolved, after which they elicit fewer notable actions.
Homeostatic Intrinsic Motivation
Recently, studies have begun to emerge that address homeostatic intrinsic motivation, in which agents autonomously define their own setpoints but aim to preserve already learned states or past experiences, rather than continuously seeking novelty. These studies have made important contributions by demonstrating that goal-directed behaviors can emerge without external rewards. For example, (marcel_learning_2022) proposed a process in which latent representations of self-touch observations are used as setpoints to induce reaching behavior, and (takemura_neural_2018; kim_amam_2023) demonstrated that forward and inverse models learned from motor babbling can generate reaching toward target positions without explicit rewards. Most similar to our approach, (sajid_exploration_2021) introduced the concept of "learned preferences" within the free energy framework and semantically analyzed how a drive to reduce mismatches between observation and preference can lead to new behaviors. While these studies have demonstrated the potential of homeostatic intrinsic motivation, they operate within independent frameworks focused on a single intrinsic motivation mechanism, and integration with extrinsic motivations remains a future challenge.
Summary
In summary, our approach is differentiated from existing research in the following ways:
  1. 1.
    Intrinsic: The agent autonomously establishes criteria based on its own experience rather than adhering to externally fixed criteria.
  2. 2.
    Independent behavior generation: It independently generates specific goal-directed behaviors such as reaching, rather than serving an auxiliary role to facilitate exploration.
  3. 3.
    Homeostatic: Even in situations without information gain, behaviors persist without stopping as long as there is a mismatch with familiar states.
  4. 4.
    FEP with active inference: Defined within the active inference framework, it can be theoretically integrated with existing extrinsic motivations and heterostatic intrinsic motivations (information gain).
@@ 5.2 Origin of Early Intentional Behaviors
Infants initially depend heavily on reflex mechanisms for action. Over time, however, they gradually develop what (mele_intentional_1994) describes as “intentional behavior,” comprising (i) a motivational system that judges how to handle incoming stimuli and (ii) the capability to execute genuinely goal-directed actions. According to (zaadnoordijk_origins_2020), stimulus-driven intentions arise first, and then, with accumulating experience, more endogenous intentions emerge. For instance, newborns rely on reflexes such as rooting to acquire feeding experience; later, merely seeing a bottle prompts them to reach out and drink (stimulus-driven intention). As they become aware of internal states like hunger, they begin actively searching for the bottle themselves (endogenous intention).
@@
Over time, to recognize themselves as independent entities, infants must undergo abundant experiences featuring multimodal redundancy and temporal–spatial contingencies (rochat_self-perception_1998). Sequentially developing intentional actions may spontaneously provide opportunities for such experiences. (thelen_transition_1993) similarly noted that infants initially lack both the ability and the intention to perform reaching, and that the formation of stable intentions, coupled with active exploration driven by those intentions, facilitates the learning of reaching.
However, prior studies have not discussed how intentions, including stimulus-driven intentions, and intentional behaviors develop through what structure. Even (priorelli_flexible_2023), which implemented switchable intentions within the active inference framework, used intentions that were discretely predefined rather than self-developing. Meanwhile, neurophysiological evidence suggests that the anterior cingulate cortex primarily deals with reward-related errors in individual decisions, whereas the posterior cingulate cortex processes discrepancy signals related to autobiographical information such as one’s own body or past experiences, and triggers new behaviors (pearson_posterior_2011; brewer_what_2013).
@@
Our proposed self-prior aligns with this notion of “self-relevant predictions,” indicating that when new input fails to satisfy these predictions, the system initiates action. In other words, even without any externally designed goal, an intrinsic drive to preserve a self-prior learned from past experiences can manifest behaviors akin to an infant persistently touching a newly discovered sticker on their arm. This mirrors Piaget’s concept of primary circular reactions, where repeated attempts can refine intentions further, offering a theoretical foundation for computational studies on the origins of infant intentional behavior.
@@
In our experiments, we fixed the self-prior’s learning rate. However, an excessively high rate could lead to overreactions (compulsive-like responses) to minor changes, whereas an overly low rate might cause indifference (apathy) toward new stimuli. Future work could investigate how more flexible learning rates shape different behavioral styles via self-relevant prediction-error processing.
@@ 5.3 Limitations and Future Directions
Our study employs a constructivist approach, focusing on demonstrating how the self-prior mechanism operates under minimal conditions. The infant sticker-reaching task serves as one example to clearly illustrate this mechanism, showing what behaviors emerge under the specific condition of tactile and proprioceptive body reaching. Since the self-prior is modality-independent in principle, the same mechanism can be applied to various sensory modalities. For example, using visual modality would be expected to produce reaching behavior toward rarely seen external objects outside the agent’s body. Indeed, several phenomena observed in actual infant development may also be explained by the self-prior mechanism. For instance, according to (eskandari_effect_2020), preterm infants nested in boundaries showed a significant reduction in unstable extension movements and maintained stable flexed postures, which can be reinterpreted as the infants preferring uterine-like environments and maintaining stable behavioral states to resolve mismatches with a self-prior learned from the flexed posture and tactile boundary sensations experienced during fetal stages.
Furthermore, behaviors that emerge in more complex cognitive contexts beyond simple sensorimotor-level adjustments, such as attention or social interactions, may also be explained by the self-prior. (warneken_altruistic_2006) presented tasks in which infants help adults who are having difficulty achieving goals, including results where infants opened a door for an adult who was having difficulty putting a magazine into a closet. Reinterpreting this in terms of the self-prior approach, it is possible that the infant opened the door to resolve a mismatch with a self-prior learned from visual information that adults normally open the closet door and put magazines in. Thus, the same principle of learning the probability density over experienced observations can explain behaviors across diverse contexts. Our approach has only been validated in discrete environments and low-degree-of-freedom (3-DoF) continuous environments, but real infants or robots exhibit far more complex and varied sensorimotor interactions. Therefore, it is necessary to extend this study to high-dimensional state and action environments. In particular, by utilizing simulations of early human sensorimotor experiences including fetal stages (kim_simulating_2022), it would be possible to explore how rich sensorimotor inputs provided in utero shape the self-prior and goal-directed behaviors, thereby revealing what behaviors can develop under physical constraints (kuniyoshi_fusing_2019).
In such complex environments, achieving generalizations such as reusing priors acquired across diverse contexts would require integrating all available modalities, learning various scenes over extended periods, and employing high-dimensional self-prior models capable of probabilistically modeling all of these. Such models could potentially be achieved through hierarchical active inference (friston_graphical_2017), or deep generative models like Transformers for handling long time series (vaswani_transformer_2017; chen_transdreamer_2022). Since the self-prior learns experienced sensations in a density model without prior knowledge, such extensions are possible in principle through scaling up generative models.
For example, generalizations such as transferring reaching behavior learned through touch to visual stimuli, generalizing responses to a specific shape of sticker to other shapes, or extending behaviors toward objects attached to one’s own body to external objects are all expected to be possible through large-scale generative models capable of learning statistical regularities across diverse sensory experiences. Such generalization can be understood not simply as repetition of the same action, but as a process of recognizing and utilizing structural similarities among self-priors formed in different contexts. Additionally, we showed behaviors driven solely by familiarity based intrinsic motivation by excluding externally imposed reward criteria, but in reality, infants or robots must balance multiple motivations such as energy homeostasis and novel information exploration. Therefore, designing experiments that systematically assess how these motivations complement or conflict and organize behavior in large-scale tasks will be an important task for future research.
@@
Meanwhile, our study demonstrated that, by separating an intrinsic homeostatic term using a learnable self-prior from the extrinsic term using a fixed preferred prior that was commonly provided in prior studies using expected free energy, an agent can not only generate actions aimed at achieving predefined goals but also spontaneously generate goal-directed behaviors. However, a current limitation is that while the use of the self-prior is integrated into policy generation via expected free energy, the learning of the self-prior is performed separately from the learning of variational free energy.
@@
One potential clue to resolving this separation may also lie in a hierarchical architecture. Specifically, a higher-level hidden state that reflects long-term experiences may be gradually learned, and the likelihood distribution inferred from it could serve as the self-prior for the lower level. This differs from conventional hierarchical inference, where the higher level simply learns priors and determines actions on a slower timescale to influence the lower level (e.g., $P(s_{t}|s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$). Instead, it continuously receives information from the lower level to learn a prior, serving only to form a self-prior over a long duration (e.g., $P(s)=\text{Cat}(\mathbf{B})$). This hierarchical approach is also expected to be closely connected to building preferences over internal states. While the current implementation learns the self-prior in observation space, through a hierarchical approach, there is a possibility that higher-level internal latent representations may be formed through experience.
@@
Finally, in the discrete environment, we could observe that as time passed with a sticker attached, the self-prior gradually changed and, accordingly, behavior also changed. By contrast, in the continuous environment, we conducted the experiment by fixing the probability of stickers being included in the observations used for self-prior learning at 5%. This was because the self-prior model stored past data in a replay buffer and learned without considering temporal order, making it difficult to investigate gradual changes in the self-prior according to changes in observation frequency. To reproduce the process by which stimulus-driven intention gradually strengthens into a more endogenous form of intention in real infants or robots, a self-prior learning mechanism that considers temporal order is necessary. This could potentially be implemented through the hierarchical architecture mentioned earlier, where higher levels integrate long-term experiences.